OpenStudy (anonymous):

solve linear equation by graphing 3x+y=0 2x+y=-3

OpenStudy (zzr0ck3r):

do you know how to graph linear functions?

OpenStudy (anonymous):

nope

OpenStudy (zzr0ck3r):

y = -3x for the first y = -2x-3

OpenStudy (zzr0ck3r):

do you know what a slop and a y intercept is?

OpenStudy (zzr0ck3r):

slope*

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

do you know what each variable in y = m+b is for?

OpenStudy (zzr0ck3r):

err y = mx+b

OpenStudy (zzr0ck3r):

in other words what is the slope of y = -3x ?

OpenStudy (zzr0ck3r):

@aea2213 ?

OpenStudy (anonymous):

is the slope 0

OpenStudy (zzr0ck3r):

no y = mx+b m = slope b = y intercept so for y = -3x we have no b value, so the y intercept is 0, and the slope is -3 make sense?

OpenStudy (anonymous):

yea so the pair would be (0,-3)

OpenStudy (zzr0ck3r):

nono, -3 is a slope, it is not a y value. so we would draw a graph, and start at the y intercept, wich is (0,0) and then go down 3 to the right 1 and make a point, then go down 3 and to the right 1....... make sense?

OpenStudy (zzr0ck3r):

do you understand why the y intercept is (0,0)?

OpenStudy (anonymous):

ok kinda i understand

OpenStudy (anonymous):

because there is no y intercept

OpenStudy (zzr0ck3r):

ok what is the slope and y intercept (ordered pair) of y = -2x-3

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

2 is the slope and -3 y intercept

OpenStudy (zzr0ck3r):

well there is a y intercept, but since there is no b value, b must be = to 0

OpenStudy (zzr0ck3r):

slope is -2

OpenStudy (zzr0ck3r):

but wait back up a sec.

OpenStudy (zzr0ck3r):

on the last one y = -3x is the same thing as y = -3x+0 thus our y intercept is 0 and when we talk about a y intercept we are saying x = 0 so its the ordered pair (0,0) ok?

OpenStudy (zzr0ck3r):

now for y = -2x-3 the slope is -2 and the y intercept is -3

OpenStudy (zzr0ck3r):

so we know the graph crosses the y axis when x = 0 and y = -3 or at (0,-3)

OpenStudy (zzr0ck3r):

do you see why the slope is -2 and n ot 2? we always take the sign of m

OpenStudy (anonymous):

no i dont :/

OpenStudy (anonymous):

but so I graph (0,-3) and (-2,-3)

OpenStudy (zzr0ck3r):

um you start at the y intercept (0,-3) and then since the slope is -2 and -2 = -2/1 and slope is rise/run we go down 2 then to the right 1. then down 2 and to the right 1....

OpenStudy (anonymous):

ok thanks